On January 1, Graves Corporation had 60,000 shares of no-par common stock issued and outstanding. The stock has a stated value of $4 per share. During the year, the following transactions occurred. Apr. 1 Issued 9,000 additional shares of common stock for $11 per share. June 15 Declared a cash dividend of $1.50 per share to stockholders of record on June 30. July 10 Paid the $1.50 cash dividend. Dec. 1 Issued 4,000 additional shares of common stock for $12 per share. Dec. 15 Declared a cash dividend on outstanding shares of $1.60 per share to stockholders of record on December 31.

(b) How are dividends and dividends payable reported in the financial statements prepared at December 31? type your answer here:
Answer #1

Dividends declared on June 15 = (60,000 + 9,000) * $1.50 = $103,500

Dividends declared on Dec 15 = (60,000 + 9,000 + 4,000) * $1.60 = $116,800

Total dividends declared during the year = $103,500 + $116,800 = $220,300

Dividends declared on Dec 15 are still payable on December 31

Reporting in financial statements prepared at December 31 shall be as follows -

i) Dividends of $220,300 shall be shown as a negative item in the Statement of Retained Earnings.

ii) Dividends payable of $116,800 shall be shown as current liabilities in the Balance Sheet.